solve the equation n^2+6n+4=0 by compleating the square
Answer by nerdybill(7384)

n^2+6n+4=0
n^2+6n = -4
take 1/2 b-coefficient, square it, and add it to both sides:
n^2+6n+9 = -4+9
(n+3)(n+3) = 5
(n+3)^2 = 5
n+3 = sqrt(5)
n = +-sqrt(5) - 3
.
n = {-0.764, -5.236}